solve -11<4x-4x<-3
the inequaility signs are less then equal to signs

Answer by RAY100(1637)

As it stands the 4x-4x=0,,,and this is not a reasonable problem.
.
If your computer is like mine,,,,this is probably,,,
.
-11<= 4x-4 <= -3,,,,,,if so ok,,,if not pls resub problem
.
lets start by adding 4 to all terms
.
-11+4 <= 4x -4+4 <+ -3 +4
.
-7 <= 4x <= 1
.
lets now divide by 4
.
-7/4 <= 4x/4 <= 1/4
.
-1.75 <= x <= .25
.
.
check,,, let x= -1,,,(-1.75<= x),,,,-11<= -8<= -3
.
let x=.2,,(x<=.25),,,,-11<=-3.2 <= -3,,,,ok